OpenStudy (pphalke):

how many liters of a 60% acid solution must be mixed with a 30% acid solution to get 300L of a 50% acid solution?

OpenStudy (snowsurf):

These mixture problems are easy when you create a table. Let's do that. Let x = 30% acid solution and y = 60% acid solution 

OpenStudy (snowsurf):

So from this table you add up the total liters which is x + y = 300 And the total solution is 0.30x + 0.60y = 150 In the table I should had put the total liters as 300 L instead of x + y because x + y that sum equations a total of 300 liters. So using those two equations you can find x and y. You can use substitution of elimination method to find the solution.

OpenStudy (pphalke):

how do i solve the equation

OpenStudy (pphalke):

would it be 0.30x+0.60y=150 or 300?

OpenStudy (snowsurf):

Let's do substitution method. x + y = 300. Let's use this equation. Solve for y. It doesn't have to be y it could b x but I just choose y for this purpose. y = 300 - x Then you take that equation plug it into this one 0.30x + 0.60y = 150 So you get 0.30x + 0.60 (300 - x) = 150. Solve for x and once you have x then you plug into either equation 0.30x + 0.60y = 150 or x + y = 300 to find y.

OpenStudy (pphalke):

i dont know how to find y

OpenStudy (snowsurf):

Plug x into this equation y = 300 - x

OpenStudy (pphalke):

so it would be y=300-0.30?

OpenStudy (snowsurf):

No you have to solve this equation 0.30x + 0.60 (300 - x) = 150 to find x. What step did I do you find most confusing?

OpenStudy (pphalke):

how do i get the answer ? im still confused

OpenStudy (snowsurf):

1.) x + y = 300 2.) 0.30x + 0.60y = 150 So you solve for one of the variables. Because you cannot solve an equation with two unknowns. So equation 1.) I will solve for y. When I do equation 1.) becomes y = 300 - x Then I take that equation plug it into this equation 0.30x + 0.60y = 150. When I do that I get the following 0.30x + 0.60 (300 - x) = 150 Solve for x. You use your algebra. So you distribute the 0.60 (300 - x ) = 0.60*300 - 0.60x 0.30x + 180 - 0.60x = 150 Combine like terms -0.30x + 180 = 150 subtract 180 on both sides -0.30x = -30 divide -0.30 on both sides to solve for x x = 100 Now you know x = 100 plug x into the equation x + y = 300 to find y

OpenStudy (pphalke):

is it 100+y=300?

OpenStudy (snowsurf):

Yes. Now solve for y and you will have your solution.

OpenStudy (pphalke):

can you write it like this y+100=300?

OpenStudy (snowsurf):

Yeah you can if you want. The idea is to get y by itself 100+y=300 You subtract 100 on both sides 100 - 100 + y = 300 - 100 y = 200

OpenStudy (pphalke):

is that the answer to the question??

OpenStudy (snowsurf):

yes. x = 100 and y = 200

OpenStudy (pphalke):

so 200 liters of a 60% acid solution must be used?

OpenStudy (snowsurf):

That is correct